Suppose there are 10 liters of a gas at 20 degrees Celsius. If the pressure does not change, what will be the volume of the gas at O degree Celsius


1
Expert's answer
2021-12-17T03:10:02-0500

"\frac{V_1}{T_1}=
\frac{V_2}{T_2}"


"V_2=\frac{10×273}{293}=9.32litres"